If Joshua paid for 48,000$ for all of his expenses, consiting of dresses and shoes, with each dress being 400$ and each pair of shoes being 1000$, after he sold all of then he had a profit of 18,000$ with a profit of 20% for the dresses, and profit of 50% for the shoes, how many shoes and dresses did Joshua buy?
Answer by Theo(13342)

let x = number of dresses and y = number of shoes.

the cost of each dress is 400 dollars and the cost of each pair of shoes is 1000 dollars.

the total cost is 48,000.

therefore:

400x + 1000y = 48,000.

that equation says that the number of dresses * 400 plus the number of pairs of shoes * 1000 equals the total cost of 48000.

the total profit is 18000.

the profit margin on dresses is 20%.
the profit margin on shoes is 50%.

the profit on the dresses is 20% of the cost of the dresses.

therefore, the profit on the dresses is .20 * 400 * x.

the profit on the shoes is 50% of the cost of the shoes.

therefore, the profit on the shoes is .50 * 1000 * y.

since the total profit is 18,000, you get:

.20 * 400 * x + .50 * 1000 * y = 18,000.

looks like we have 2 equations that need to be solved simultaneously.

they are:

400x + 1000y = 48,000
.20 * 400 * x + .50 * 1000 * y = 18,000

simplify the second equation and leave the first equation as is to get:

400x + 1000y = 48,000
80x + 500y = 18,000

leave the first equation as is and multiply both sides of the second equation by 2 to get:

400x + 1000y = 48,000
160x + 1000y = 36,000

subtract the second equation from the first to get:

240x = 12,000

solve for x to get:

x = 12000 / 240 = 50

you have x = 50

your 2 original equations, after simplification, are:

400x + 1000y = 48,000
80x + 500y = 18,000

replace x with 50 in the first original equation to get:

400*50 + 1000y = 48,000

simplify to get:

20,000 + 1000y = 48,000

solve for y to get y = (48,000 - 20,000) / 1000 = 28.

you have x = 50 and y = 28

replace x with 50 and y with 28 in the second original equation to get:

80x + 500y = 18,000 becomes:

80 * 50 + 500 * 28 = 18,000

simplify to get 4,000 + 14,000 = 18,000.

both original equations are true when x = 50 and y = 28, so that must be your solution.

she bought 50 dresses and 28 pairs of shoes.

the dresses cost her 50 * 400 = 20,000.
the shoes cost her 28 * 1000 = 28,000.

her total cost was 20,000 + 28,000 = 48,000.

she made 20% profit on the dresses.
she made 50% profit on the shows.

her profit on the dresses was .20 * 20,000 = 4,000.
her profit on the shoes was .50 * 28,000 = 14,000.

her total profit was 4,000 + 14,000 = 18,000.

solution checks out and is therefore assumed to be good.

solution is 50 dresses and 28 pairs of shoes were bought.